OpenStudy (anonymous):

4r*2r+πr^2 factored form?

OpenStudy (zzr0ck3r):

\(4r2r+\pi r^2=8r^2+\pi r^2 = r^2(8+\pi)\)

OpenStudy (anonymous):

thanks